# Bridge multisections of knotted surfaces in $S^{4}$

### v1

@@ 4. Band surgery on multiplane diagrams
In [MTZ23], the authors described local modifications for triplane diagrams that affect the embedded surface by taking the connected sum with an unknotted $RP^{2}$ and by performing 1–handle addition. In this section, we explore these operations in the context of bridge multisections. We have three applications of these moves: Theorem 7.2 shows that any $n$–regular Tait graph can appear as the spine of a bridge multisection for an unknotted surface, Theorem 5.8 is a uniqueness result for multiplane diagrams yielding the same surface, and Theorem 6.3 is a uniqueness result for all multiplane diagrams.
@@ 4.1. Neighborhoods of a point
One can build new multiplane diagrams from old ones by performing local modifications around points of the central surface $\Sigma$. For instance, a multiplane diagram for the connected sum of two multisected surfaces can be obtained by taking the connected sum of each pair of tangles along the same puncture (see Section 3.3). Similarly, the elementary perturbation of a triplane diagram involves replacing a certain neighborhood of an interior point of the central surface $*\in int(\Sigma)$ by a suitable properly embedded surface in the 4-ball. Below we describe other ways in which a 4-ball neighborhood of $*\in int(\Sigma)$ can interact with a given multisected surface.
The goal is to describe 1–handle additions and crosscap sums using framed arcs in each tangle of a multiplane diagram. As described in Definition 4.1, we study two kinds of arcs: those that intersect a tangle (type 1) and those that don’t (type 0). Lemma 4.2 explains how the arc data determines a disk neighborhood of $*\in int(\Sigma)$. This is important as we want to understand what part of the multisected surface we are modifying.
@@
Let $(S^{4},F)=\bigcup_{i=1}^{n}(X_{i},\mathcal{D}_{i})$ be a bridge multisection with multiplane diagram $\mathcal{T}$ and let $*\in int(\Sigma)$ be an interior point of the central surface. A collection of arcs $\rho=\{\rho_{i}\}_{i=1}^{n}$ is called admissible if it satisfies the following properties:
@@
    if both $\rho_{i}$ and $\rho_{i+1}$ are type 1 and $\rho_{i}\cup\rho_{i+1}$ connects distinct components $l$ and $l^{\prime}$ of $L_{i}$, then there are pairwise disjoint embedded 2–spheres $S$ and $S^{\prime}$ in $B_{i}\cup\overline{B}_{i+1}$ satisfying that
@@
      $S\cap\Sigma$ and $S^{\prime}\cap\Sigma$ are loops in $\Sigma$ not containing *.
Conditions (3) and (4) above may appear a bit cumbersome. In practice, one can interpret them as follows: after some isotopies of $L_{i}\cup\rho_{i}\cup\rho_{i+1}$ fixing the bridge surface $\Sigma$, $L_{i}\cup\rho_{i}\cup\rho_{i+1}$ has a diagram for which $\alpha\cup\rho_{i}\cup\rho_{i+1}$ (resp. $l\cup l^{\prime}\cup\rho_{i}\cup\rho_{i+1}$) is crossingless like in the left (resp. right) side of Figure 7. Figures 1 and 10 have arcs satisfying Conditions (3) and (4), respectively. Another feature of these conditions is that one can find embedded disks $D_{i}\subset\partial X_{i}$ bounded by $L_{i}$ and interior disjoint from $\rho_{i}\cup\rho_{i+1}$. If the arcs satisfy condition (3), the disks $D_{i}$ can be chosen so that $\rho_{i}\cup\rho_{i+1}$ is isotopic to an arc in $D_{i}$.
@@
Let $\rho$ be a collection of admissible arcs for $\mathcal{T}$. For every $i=1,\dots n$, there is an embedded disk $\Delta_{i}\subset X_{i}$ with interior in $int(X_{i})$ satisfying the following conditions:
@@
      In the other cases [^1] Some of $\rho_{i}$, $\rho_{i+1}$ is type 0 or both are type 1 with $\rho_{i}\cup\rho_{i+1}$ connecting distinct components of $L_{i}$, $\delta_{i}$ is an arc with interior disjoint from $\mathcal{D}_{i}$.
@@
Uniqueness of $\mathcal{D}_{i}$ is the statement of Lemma 2.2. Conditions (1) and (2) for $\Delta_{i}$ above imply that the disk $\Delta_{i}$, depicted as a properly embedded disk in the complement of $\mathcal{D}_{i}$, is a trivial disk system. As $\overline{B^{4}-N(\mathcal{D}_{i})}$ is diffeomorphic to a connected sum of copies of $S^{1}\times B^{3}$, Lemma 8 of [MZ18] implies that $\Delta_{i}$ is the unique disk system up to isotopy fixing $\delta\cup\rho_{1}\cup\rho_{2}$. Thus, any other disk $\Delta^{\prime}_{i}$ for $\mathcal{D}_{i}$ must be isotopic to $\Delta_{i}$ (rel boundary) in the complement of $\mathcal{D}_{i}$.
To prove existence, we have cases to discuss, depending on the types of $\rho_{i}$ and $\rho_{i+1}$ and if $\rho_{i}\cup\rho_{i+1}$ connects the same component of $L_{i}$ or not. If one of the arcs is type 0, say $\rho_{i}$, we can shrink it so it lies near the point * and then enlarge it so that it is equal to $\rho_{i+1}$. This movie traces the desired disk $\Delta_{i}$.
@@
    If $\rho_{i}\cup\rho_{i+1}$ connects the same component of $L_{i}$, then there exist exactly two framings, called the ±$\pm 1$–framings $f$, such that $L_{i}[\rho_{i}\cup\rho_{i+1},f]$ is a $c_{i}$–component unlink. Moreover, there exist exactly one framing, called the zero-framing $f$, such that $L_{i}[\rho_{i}\cup\rho_{i+1},f]$ is a $(c_{i}+1)$–component unlink.
@@
Conditions (3) and (4) of Definition 4.1 ensure in each case the existence of disks $D_{i}\subset\partial X_{i}$ with boundary $L_{i}$ and interior disjoint from $\rho_{i}\cup\rho_{i+1}$. If $\rho_{i}\cup\rho_{i+1}$ connects the same component of $L_{i}$, such disks imply that $L_{i}[\rho_{i}\cup\rho_{i+1},f]$ is an unlink of $(c_{i}-2)$ unknots together with a $(2,n)$–torus link, where $n$ depends on the framing of $\rho_{i}\cup\rho_{i+1}$. In this case, there are exactly two framings ($n=\pm 1$) for which the torus knot is an unknot and one framing ($n=0$) for which we have a 2–component unlink. On the other hand, if $\rho_{i}\cup\rho_{i+1}$ connects distinct components of $L_{i}$, the band given by any framing $f$ becomes a connected sum band. Hence, $L_{i}[\rho_{i}\cup\rho_{i+1},f]$ is a $(c_{i}-1)$–component unlink. ∎
@@
Let $(\rho,f)=\{(\rho_{i},f_{i})\}_{i=1}^{n}$ be a tuple of framed arcs in a multiplane diagram $\mathcal{T}$. We say that $(\rho,f)$ is admissible if $\rho$ is a collection of admissible arcs for $\mathcal{T}$, the framings $\{f_{i}\}_{i=1}^{n}$ agree at $*\in int(\Sigma)$, and the framing $f_{i}+f_{i+1}$ is zero or ±$\pm 1$ whenever $\rho_{i}\cup\rho_{i+1}$ connects the same component of $L_{i}$.
@@ 4.2. Band surgery
A multiplane diagram $\mathcal{T}=(T_{1},\dots,T_{n})$ can be surgered along an admissible tuple of framed arcs $(\rho,f)$ to yield a new tuple of trivial tangles denoted by $\mathcal{T}[\rho,f]=(T_{1}[\rho_{1},f_{1}],\dots,T_{n}[\rho_{n},f_{n}])$. We say that $\mathcal{T}[\rho,f]$ is the result of performing band surgery to $\mathcal{T}$ along $(\rho,f)$. We call this operation band surgery as bands (thus surgered tangles) are obtained from framed arcs as in Remark 2.3.
@@
If $(\rho,f)$ is an admissible tuple of framed arcs for a $b$–bridge multiplane diagram $\mathcal{T}$, then $\mathcal{T}[\rho,f]$ is the spine of a $(b+1)$–bridge multisected surface.
@@
We need to check that the surgered link $L_{i}[\rho_{i}\cup\rho_{i+1},f_{i}+f_{i+1}]$ is an unlink in $(b+1)$–bridge position for each pair of consecutive arcs $\rho_{i}$ and $\rho_{i+1}$. Lemma 4.3 ensures this whenever $\rho_{i}$ and $\rho_{i+1}$ are type 1. If both arcs are type zero, then $L_{i}[\rho_{i}\cup\rho_{i+1},f_{i}+f_{i+1}]$ is obtained from $L_{i}$ by adding a small 1–bridge unknot. If $\rho_{i}$ and $\rho_{i+1}$ are of different types, then $L_{i}[\rho_{i}\cup\rho_{i+1},f_{i}+f_{i+1}]$ is obtained by perturbing $L_{i}$. ∎
@@
Consider an admissible tuple $(\rho,f)$ and disks $\{\Delta_{i}\}_{i=1}^{n}$ as in Lemma 4.2. By construction, the surface represented by $\mathcal{T}[\rho,f]$ differs from $F$ in a neighborhood of the disks $\Delta_{1}\cup\dots\cup\Delta_{n}$. The new disk system $\mathcal{D}^{\prime}_{i}$ is obtained from $\mathcal{D}_{i}$ by replacing a neighborhood of $\Delta_{i}\cap\mathcal{D}_{i}$ with certain push-offs of $\Delta_{i}$. Figure 9 contains drawings of these modifications for the different types of $\rho_{i}$ and $\rho_{i+1}$ which we now describe. If both arcs $\rho_{i}$ and $\rho_{i+1}$ are type 0, then $\Delta_{i}$ is disjoint from $\mathcal{D}_{i}$. In this case, $\mathcal{D}^{\prime}_{i}$ gains a small disk near $\Delta_{i}$ as in Part (a) of Figure 9. If the types of $\rho_{i}$ and $\rho_{i+1}$ are distinct as in Part (b), the disk $\mathcal{D}^{\prime}_{i}$ is an elongated copy of $\mathcal{D}_{i}$. When both $\rho_{i}$ and $\rho_{i+1}$ are type 1 and $\rho_{i}\cup\rho_{i+1}$ connects the same component of $L_{i}$, $\mathcal{D}_{i}$ intersects a neighborhood of $\Delta_{i}$ in a rectangle as in Part (c). Here, if the framing $f_{i}+f_{i+1}$ is zero, then we replace this rectangle with two disjoint copies of $\Delta_{i}$. If the framing is ±$\pm 1$ instead, we add a half-twisted band connecting the two copies of $\Delta_{i}$ as in the right side of Part (c). As we see in Part (d), if $\rho_{i}+\rho_{i+1}$ connects distinct components of $L_{i}$, $\mathcal{D}_{i}$ gains a band with twisting given by the framing of the arcs.
Figure 9. How the disks $\mathcal{D}_{i}$ change after surgery along an admissible tuple $(\rho,f)$.
@@
Before we state the main technical result of this section (Thm 4.9), we will describe examples of tuples $(\rho,f)$ for which we understand how $F_{\mathcal{T}}$ and $F_{\mathcal{T}[\rho,f]}$ are related.
@@
Fix $21\leq m\leq n-$ 2 and consider an admissible tuple $(\rho,f)$ having the following properties: (1) the arcs $\rho_{i}$ are type 1 for $i=1,\dots,m+1$ and type 0 otherwise, and (2) for $i=1,\dots,m$, the framing $f_{i}+f_{i+1}$ is zero whenever $\rho_{i}\cup\rho_{i+1}$ connects the same component of $L_{i}$. If $r$ is the number of arcs $\rho_{i}\cup\rho_{i+1}$ connecting distinct components of $L_{i}$, then the surface $F_{\mathcal{T}[\rho,f]}$ is obtained from $F_{\mathcal{T}}$ by adding $r$ 1–handles.
This fact follows from Remark 4.10. When $(n,m,r)=(n,m,0)$, $\mathcal{T}[\rho,f]$ and $\mathcal{T}$ represent isotopic surfaces. We will discuss more on this case in Section 5.1. When $(n,m,r)=(3,1,1)$ this corresponds to the tubing operation introduced in [MTZ23]. Figure 10 shows the fact that the spin of a 2–bridge knot is obtained by adding one 1–handle to a 2–component unlink of 2–spheres.
Figure 10. Turning a $(3;2,2,3)$–bridge trisection of a 2–component unlink of 2–spheres into a $(4;2)$–bridge trisection of a spun knot. The top surface is unknotted by Proposition 3.9.
@@
Consider an admissible tuple $(\rho,f)$ satisfying the following condition: (1) if $\rho_{i}$ and $\rho_{j}$ type 1 arcs for some $i\neq j$, then $12<|i-j|<n-$ 1. If $t$ is the number of type 1 arcs, then the surface $F_{\mathcal{T}[\rho,f]}$ is obtained from $F_{\mathcal{T}}$ by adding $(t-1)$ 1–handles.
This fact follows from Remark 4.10. When $t=1$, $\mathcal{T}[\rho,f]$ and $\mathcal{T}$ represent isotopic surfaces (see Remark 5.3 for more on this case). Figure 2 shows how to use one 1–handle to turn a 2–component unlink of 2–spheres into the spin of a trefoil knot. We leave it as an exercise for the reader to see that this multiplane diagram is equivalent to some $\mathcal{T}(\beta,(0,0))$ from Proposition 3.11.
@@
Fix $21\leq m\leq n-$ 2 and consider an admissible tuple $(\rho,f)$ having the following properties: (1) the arcs $\rho_{i}$ are type 1 for $i=1,\dots,m+1$ and type 0 otherwise, and (2) for $i=1,\dots,m$, the arc $\rho_{i}\cup\rho_{i+1}$ connects the same component of $L_{i}$. If $r$ is the number of framings $f_{i}+f_{i+1}$ equal to ±$\pm 1$ for some $i=1,\dots,m$. Then the surface $F_{\mathcal{T}[\rho,f]}$ is obtained from $F_{\mathcal{T}}$ by taking the connected sum with $r$ unknotted projective planes.
This fact follows from Remark 4.10. When $(n,m,r)=(3,1,1)$, this corresponds to the crosscap sum introduced in [MTZ23]. In Figure 1, we show how to draw a 4-plane diagram for an unknotted Klein bottle by adding two projective planes to a 1–bridge diagram of a 2–sphere.
The theorem below shows that most band surgeries change the underlying surface by 1–handle additions and crosscap sums. In practice, Remark 4.10 gives a concrete process to decide how many of which tubes and crosscaps are needed for a given admissible tuple.

### v2

@@ 4. Band surgery on multiplane diagrams
In [MTZ23], the authors described local modifications for triplane diagrams that affect the embedded surface by taking the connected sum with an unknotted $RP^{2}$ and by performing 1–handle addition. In this section, we explore these operations in the context of bridge multisections. We have three applications of these moves: Theorem 7.2 shows that any nn–regular Tait graph can appear as the spine of a bridge multisection for an unknotted surface, Theorem 5.10 is a uniqueness result for multiplane diagrams yielding the same surface, and Theorem 6.3 is a uniqueness result for all multiplane diagrams.
@@ 4.1. Neighborhoods of a point
One can build new multiplane diagrams from old ones by performing local modifications around points of the central surface $\Sigma$. For instance, a multiplane diagram for the connected sum of two multisected surfaces can be obtained by taking the connected sum of each pair of tangles along the same puncture (see Section 3.3). Similarly, the elementary perturbation of a triplane diagram involves replacing a certain neighborhood of an interior point of the central surface $\in int(\Sigma)$ by a suitable properly embedded surface in the 4-ball. Below we describe other ways in which a 4-ball neighborhood of $\in int(\Sigma)$ can interact with a given multisected surface.
The goal is to describe 1–handle additions and crosscap sums using framed arcs in each tangle of a multiplane diagram. Definition 4.1 studies two kinds of arcs: those that intersect a tangle (type 1) and those that don’t (type 0). For this, recall the discussion on bands from Section 2.2; specifically panels (b) and (c) in Figure 4. Lemma 4.2 explains how the arc data determines a disk neighborhood of $\in int(\Sigma)$. This is important as we want to understand what part of the multisected surface we are modifying.
@@
Let $(S^{4},F)=\bigcup_{i=1}^{n}(X_{i},\mathcal{D}_{i})$ be a bridge multisection with multiplane diagram $\mathcal{T}$ and let $\in int(\Sigma)$ be an interior point of the central surface. A collection of arcs $\rho=\{\rho_{i}\}_{i=1}^{n}$ is called admissible if it satisfies the following properties:
@@
    if both $\rho_{i}$ and $\rho_{i+1}$ are type 1 and $\rho_{i}\cup\rho_{i+1}$ connects distinct components $l_{i}$ and $l_{i}^{\prime}$ of $L_{i}$, then there are pairwise disjoint embedded 2–spheres $S$ and $S^{\prime}$ in $B_{i}\cup\overline{B}_{i+1}$ satisfying that
@@
      $S\cap\Sigma$ and $S^{\prime}\cap\Sigma$ are loops containing some punctures of $\Sigma$ but not containing *.
Conditions (3) and (4) above may appear a bit cumbersome. In practice, one can interpret them as follows: after some isotopies of $L_{i}\cup\rho_{i}\cup\rho_{i+1}$ fixing the bridge surface $\Sigma$, $L_{i}\cup\rho_{i}\cup\rho_{i+1}$ has a diagram for which $\alpha\cup\rho_{i}\cup\rho_{i+1}$ (resp. $l_{i}\cup l_{i}^{\prime}\cup\rho_{i}\cup\rho_{i+1}$) is crossingless like in the left (resp. right) side of Figure 7. Figures 1 and 10 have arcs satisfying Conditions (3) and (4), respectively. Another feature of these conditions is that one can find embedded disks $D_{i}\subset\partial X_{i}$ bounded by $L_{i}$ and interior disjoint from $\rho_{i}\cup\rho_{i+1}$. If the arcs satisfy condition (3), the disks $D_{i}$ can be chosen so that $\rho_{i}\cup\rho_{i+1}$ is isotopic to an arc in $D_{i}$.
@@
Let $\rho$ be a collection of admissible arcs for $\mathcal{T}$. For every $i=1,\dots,n$, there is an embedded disk $\Delta_{i}\subset X_{i}$ with interior in $int(X_{i})$ satisfying the following conditions; see Figure 9 for a model of $\Delta_{i}$.
@@
      In the other cases [^2] Some of $\rho_{i}$, $\rho_{i+1}$ is type 0 or both are type 1 with $\rho_{i}\cup\rho_{i+1}$ connecting distinct components of $L_{i}$, $int(\delta_{i})\cap\mathcal{D}_{i}=\emptyset$ and $\Delta_{i}\cap\mathcal{D}_{i}=(\rho_{i}\cap T_{i})\cup(\rho_{i+1}\cap T_{i+1})$.
@@
Uniqueness of $\mathcal{D}_{i}$ is the statement of Lemma 2.2. Conditions (1) and (2) for $\Delta_{i}$ above imply that the disk $\Delta_{i}$, depicted as a properly embedded disk in the complement of $\mathcal{D}_{i}$, is a trivial disk system. As $\overline{B^{4}-N(\mathcal{D}_{i})}$ is diffeomorphic to a connected sum of copies of $S^{1}\times B^{3}$, Lemma 8 of [MZ18] implies that $\Delta_{i}$ is the unique disk system up to isotopy fixing $\delta_{i}\cup\rho_{i}\cup\rho_{i+1}$. Thus, any other disk $\Delta^{\prime}_{i}$ for $\mathcal{D}_{i}$ must be isotopic to $\Delta_{i}$ (rel boundary) in the complement of $\mathcal{D}_{i}$.
To prove existence, we have cases to discuss, depending on the types of $\rho_{i}$ and $\rho_{i+1}$ and if $\rho_{i}\cup\rho_{i+1}$ connects the same component of $L_{i}$ or not. If one of the arcs is type 0, say $\rho_{i}$, we can shrink it so it lies near the point * and then enlarge it so that it is equal to $\rho_{i+1}$. This process describes an isotopy of $\rho_{i}$ into $\rho_{i+1}$ that starts and ends in $\partial B^{4}$ and occurs in the interior of $B^{4}$ at the rest of the times; the desired disk $\Delta_{i}$ is the trace of the isotopy.
@@
In the following lemma, $L[\rho,f]$ denotes band surgery along an band described by a framed arc $(\rho,f)$ as in Section 2.2.
@@
    If $\rho_{i}\cup\rho_{i+1}$ connects the same component of $L_{i}$, then there exist exactly two framings, called the $\pm 1$–framings $f$, such that $L_{i}[\rho_{i}\cup\rho_{i+1},f]$ is a $c_{i}$–component unlink. Moreover, there exist exactly one framing, called the zero-framing $f$, such that $L_{i}[\rho_{i}\cup\rho_{i+1},f]$ is a $(c_{i}+1)$–component unlink.
@@
Conditions (3) and (4) of Definition 4.1 ensure in each case the existence of disks $D_{i}\subset\partial X_{i}$ with boundary $L_{i}$ and interior disjoint from $\rho_{i}\cup\rho_{i+1}$. In particular, $\rho_{i}\cup\rho_{i+1}$ intersects either one or two components of the $c_{i}$ components of $L_{i}$. If $\rho_{i}\cup\rho_{i+1}$ connects the same component of $L_{i}$, then $L_{i}[\rho_{i}\cup\rho_{i+1},f]$ is the union of an unlink of $(c_{i}-2)$ unknots together with a (2,n)(2,n)–torus link, where $n$ depends on the framing of $\rho_{i}\cup\rho_{i+1}$. In this case, there are exactly two framings ($n=\pm 1$) for which the torus knot is an unknot and one framing ($n=0$) for which we have a 2–component unlink. On the other hand, if $\rho_{i}\cup\rho_{i+1}$ connects distinct components of $L_{i}$, the band given by any framing $f$ becomes a connected sum band. Hence, $L_{i}[\rho_{i}\cup\rho_{i+1},f]$ is a $(c_{i}-1)$–component unlink. ∎
@@
Let $(\rho,f)=\{(\rho_{i},f_{i})\}_{i=1}^{n}$ be a tuple of framed arcs in a multiplane diagram $\mathcal{T}$. We say that $(\rho,f)$ is admissible if $\rho$ is a collection of admissible arcs for $\mathcal{T}$, the framings $\{f_{i}\}_{i=1}^{n}$ agree at $\in int(\Sigma)$, and the framing $f_{i}+f_{i+1}$ is zero or $\pm 1$ whenever $\rho_{i}\cup\rho_{i+1}$ connects the same component of $L_{i}$.
@@ 4.2. Band surgery
Definition 4.5.
Let $\mathcal{T}=(T_{1},\dots,T_{n})$ be a multiplane diagram and let $(\rho,f)$ be an admissible tuple of framed arcs for $\mathcal{T}$. The band surgery of $\mathcal{T}$ along $(\rho,f)$, denoted by $\mathcal{T}[\rho,f]=(T_{1}[\rho_{1},f_{1}],\dots,T_{n}[\rho_{n},f_{n}])$, is the new tuple where each tangle is obtained by band surgery as in Remark 2.4.
@@
If $(\rho,f)$ is an admissible tuple of framed arcs for a bb–bridge multiplane diagram $\mathcal{T}$, then $\mathcal{T}[\rho,f]$ is the spine of a (b+1)(b+1)–bridge multisected surface.
@@
We need to check that for each $i\in\mathbb{Z}_{n}$, $T[\rho_{i},f_{i}]$ is a $(b+1)$-bridge trivial tangle and the surgered link $L_{i}[\rho_{i}\cup\rho_{i+1},f_{i}+f_{i+1}]$ is an unlink. To see the first conclusion, recall that each $\rho_{i}$ is an increasing arc with one endpoint in $\in\Sigma$; see Definition 4.1(1). If $\rho_{i}$ is type 0, then $T_{i}[\rho_{i},f_{i}]$ is equal to $T_{i}$ with an added 1–bridge strand around the band described by $(\rho_{i},f_{i})$. If $\rho_{i}$ is type 1, denote by $s$ the strand of $T_{i}$ containing one endpoint of $\rho_{i}$. We can slide the arc along $s$ until we reach the unique local maximum of the strand; this is possible as $T_{i}$ is a trivial tangle. The resulting arc $\rho_{i}$ can also be assume to satisfy Condition (1) in Definition 4.1. In this position, it is easy to see that surgery along $(\rho_{i},f)$ splits $s$ into two strands that have exactly one local maximum (near the maximum of $s$); note that the two strands will twist according to the framing $f_{i}$. Thus, $T_{i}[\rho_{i},f_{i}]$ is trivial.
@@
We now explain why $L_{i}[\rho_{i}\cup\rho_{i+1},f_{i}+f_{i+1}]$ is an unlink. Lemma 4.3 ensures this whenever $\rho_{i}$ and $\rho_{i+1}$ are type 1. If both arcs are type zero, then $L_{i}[\rho_{i}\cup\rho_{i+1},f_{i}+f_{i+1}]$ is obtained from $L_{i}$ by adding a small 1–bridge unknot. If $\rho_{i}$ and $\rho_{i+1}$ are of different types, then $L_{i}[\rho_{i}\cup\rho_{i+1},f_{i}+f_{i+1}]$ is obtained by perturbing $L_{i}$. ∎
@@
Consider an admissible tuple $(\rho,f)$ and disks $\{\Delta_{i}\}_{i=1}^{n}$ as in Lemma 4.2. By construction, the surface represented by $\mathcal{T}[\rho,f]$ differs from $F$ in a neighborhood of the disks $\Delta_{1}\cup\dots\cup\Delta_{n}$. The new disk system $\mathcal{D}^{\prime}_{i}$ is obtained from $\mathcal{D}_{i}$ by modifying a neighborhood of $\Delta_{i}\cap\mathcal{D}_{i}$ with certain surfaces near $\Delta_{i}$; if $\Delta_{i}\cap\mathcal{D}_{i}$ is empty, we will add connected components to $\mathcal{D}_{i}$. Figure 9 contains drawings of these modifications for the different types of $\rho_{i}$ and $\rho_{i+1}$, which we now describe. If both arcs $\rho_{i}$ and $\rho_{i+1}$ are type 0, then $\Delta_{i}$ is disjoint from $\mathcal{D}_{i}$. In this case, $\mathcal{D}^{\prime}_{i}$ gains a small disk near $\Delta_{i}$ as in Part (a) of Figure 9. If the types of $\rho_{i}$ and $\rho_{i+1}$ are distinct as in Part (b), the disk $\mathcal{D}^{\prime}_{i}$ is an elongated copy of $\mathcal{D}_{i}$. When both $\rho_{i}$ and $\rho_{i+1}$ are type 1 and $\rho_{i}\cup\rho_{i+1}$ connects the same component of $L_{i}$, $\mathcal{D}_{i}$ intersects a neighborhood of $\Delta_{i}$ in a rectangle as in Part (c). Here, if the framing $f_{i}+f_{i+1}$ is zero, then we replace this rectangle with two disjoint copies of $\Delta_{i}$. If the framing is $\pm 1$ instead, we add a half-twisted band connecting the two copies of $\Delta_{i}$ as in the right side of Part (c). As we see in Part (d), if $\rho_{i}+\rho_{i+1}$ connects distinct components of $L_{i}$, $\mathcal{D}_{i}$ gains a band with twisting given by the framing of the arcs.
Figure 9. Given an admissible tuple $(\rho,f)$, the disks $\Delta_{i}$ can be used to change $F_{\mathcal{T}}$ in a controlled way inside each sector. We have four cases, depending on the types of the consecutive arcs $\rho_{i}$ and $\rho_{i+1}$ and whether $\rho_{i}\cup\rho_{i+1}$ connects (c) the same component of $L_{i}$ or (d) not. In case (c), we have two options, depending on the framing of $\rho_{i}\cup\rho_{i+1}$. In case (d), any choice of framing works; that said, we see that different framings produce twisting between the components of $\mathcal{D}_{i}$.
@@
Before we state the main technical result of this section (Thm 4.10), we will describe examples of tuples $(\rho,f)$ for which we understand how $F_{\mathcal{T}}$ and $F_{\mathcal{T}[\rho,f]}$ are related. We delay explanations for Examples 4.7, 4.8, and 4.9 until after Remark 4.11, as this remark provides a concrete process for determining how $F_{\mathcal{T}}$ changes.
@@
Fix $1\leq m\leq n-2$ and consider an admissible tuple $(\rho,f)$ having the following properties: (1) the arcs $\rho_{i}$ are type 1 for $i=1,\dots,m+1$ and type 0 otherwise, and (2) for $i=1,\dots,m$, the framing $f_{i}+f_{i+1}$ is zero whenever $\rho_{i}\cup\rho_{i+1}$ connects the same component of $L_{i}$. If $r$ is the number of arcs $\rho_{i}\cup\rho_{i+1}$ connecting distinct components of $L_{i}$, then the surface $F_{\mathcal{T}[\rho,f]}$ is obtained from $F_{\mathcal{T}}$ by adding $r$ 1–handles. See Example 4.12 for an explanation.
When $(n,m,r)=(n,m,0)$, $\mathcal{T}[\rho,f]$ and $\mathcal{T}$ represent isotopic surfaces. We will discuss more on this case in Section 5.1. When $(n,m,r)=(3,1,1)$ this corresponds to the tubing operation introduced in [MTZ23]. Figure 10 shows the fact that the spin of a 2–bridge knot is obtained by adding one 1–handle to a 2–component unlink of 2–spheres.
Figure 10. Turning a (3;2,2,3)(3;2,2,3)–bridge trisection of a 2–component unlink of 2–spheres into a (4;2)(4;2)–bridge trisection of a spun knot. The top surface is unknotted by Proposition 3.10.
@@
Consider an admissible tuple $(\rho,f)$ satisfying the following condition: (1) if $\rho_{i}$ and $\rho_{j}$ type 1 arcs for some $i\neq j$, then 2<|ij|<n12<|i-j|<n-1. If $t$ is the number of type 1 arcs, then the surface $F_{\mathcal{T}[\rho,f]}$ is obtained from $F_{\mathcal{T}}$ by adding (t1)(t-1) 1–handles. See Example 4.13 for an explanation.
When $t=1$, $\mathcal{T}[\rho,f]$ and $\mathcal{T}$ represent isotopic surfaces (see Remark 5.3 for more on this case). Figure 2 shows how to use one 1–handle to turn a 2–component unlink of 2–spheres into the spin of a trefoil knot. We leave it as an exercise for the reader to see that this multiplane diagram is equivalent to some $\mathcal{T}(\beta,(0,0))$ from Proposition 3.12.
@@
Fix $1\leq m\leq n-2$ and consider an admissible tuple $(\rho,f)$ having the following properties: (1) the arcs $\rho_{i}$ are type 1 for $i=1,\dots,m+1$ and type 0 otherwise, and (2) for $i=1,\dots,m$, the arc $\rho_{i}\cup\rho_{i+1}$ connects the same component of $L_{i}$. If $r$ is the number of framings $f_{i}+f_{i+1}$ equal to $\pm 1$ for some $i=1,\dots,m$. Then the surface $F_{\mathcal{T}[\rho,f]}$ is obtained from $F_{\mathcal{T}}$ by taking the connected sum with $r$ unknotted projective planes. See Example 4.14 for an explanation.
In practice, one can use Lemma 3.3 to compare the Euler numbers of $F_{\mathcal{T}}$ and $F_{\mathcal{T}[\rho,f]}$; thus determining how many crosscap summands correspond to $\mathcal{P}_{+}$ and $\mathcal{P}_{-}$. More precisely, the non-negative integers $r_{-}$ and $r_{+}$ satisfying $F_{\mathcal{T}[\rho,f]}=F_{\mathcal{T}}\#r_{-}\mathcal{P}_{-}\#r_{+}\mathcal{P}_{+}$ solve the equations
[…]
When $(n,m,r)=(3,1,1)$, this corresponds to the crosscap sum introduced in [MTZ23]. In Figure 1, we show how to draw a 4-plane diagram for an unknotted Klein bottle by adding two projective planes to a 1–bridge diagram of a 2–sphere.
The theorem below shows that most band surgeries change the underlying surface by 1–handle additions and crosscap sums. In practice, Remark 4.11 gives a concrete process to decide how many tubes and crosscaps we add to $F_{\mathcal{T}}$ to build $F_{\mathcal{T}[\rho,f]}$.